If 3x + 5x = -8, then x + 1 = ?
I can't figure out how to solve these. Please help!

Answer by Theo(13342)

you have:

3x + 5x = -8

combine like terms to get 8x = -8

solve for x to get x = -8/8 = -1

if that's true, then x + 1 = -1 + 1 = 0